OpenStudy (derpisthewordorisit):

Replace ? with =, >, or < to make the statement true. 18 ÷ 6 + 3 ? 6 + 12 ÷ 3 A. = B. < C. >

OpenStudy (derpisthewordorisit):

A I thinking.

Mehek (mehek14):

18/6=3 3+3 = 6 12/3 = 4 6 + 4 = 10 6 < 10

Mehek (mehek14):

follow PEMDAS

OpenStudy (derpisthewordorisit):

So C?